Diego Company manufactures one product that is sold for $78 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 49,000 units and sold 44,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 28
Direct labor $ 14
Variable manufacturing overhead $ 4
Variable selling and administrative $ 6
Fixed costs per year:
Fixed manufacturing overhead $ 686,000
Fixed selling and administrative expense $ 510,000

The company sold 32,000 units in the East region and 12,000 units in the West region. It determined that $230,000 of its fixed selling and administrative expense is traceable to the West region, $180,000 is traceable to the East region, and the remaining $100,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

14. Diego is considering eliminating the West region because an internally generated report suggests the region’s total gross margin in the first year of operations was $14,000 less than its traceable fixed selling and administrative expenses. Diego believes that if it drops the West region, the East region's sales will grow by 5% in Year 2. Using the contribution approach for analyzing segment profitability and assuming all else remains constant in Year 2, what would be the profit impact of dropping the West region in Year 2?

profit will____by____
Answer #1
Variable costs per unit 52 =28+14+4+6
14
Loss in Contribution margin of West region -312000 =12000*(78-52)
Fixed selling and administrative expense avoidable 230000
Increase in Contribution margin of East region 41600 =32000*5%*(78-52)
Net change in profits -40400
Profit will decrease by 40400